OpenStudy (anonymous):

what is 2tanX equal to

OpenStudy (mathmath333):

\(\large\tt \color{black}{tan2\theta=\dfrac{2tan\theta}{1-tan^2\theta}}\)

OpenStudy (acal21):

tan(2x)= 2 tan(x) / (1 - tan^2(x))---.>identity